OpenStudy (anonymous):

Given the electric potential V= -constant term-*(x^2+R^2)^1/2 - x, I was asked to find the partial derivative of V with respect to X(-dV/dx). I derived -(-constant term- *(x/((x^2+R^2)^1/2))-1) and the homework site said that was wrong. How the heck did I derive that wrong? I don't get it! Help? Excuse me if I don't seem happy. I've been doing this 5-multiple-part-problem homework assignment for the past four, almost five hours and it's frustrated me to the point of being at the verge of tears.

OpenStudy (noelgreco):

What did you get?

OpenStudy (anonymous):

I got what I said I derived.

OpenStudy (noelgreco):

Sorry. It's a lot easier for these old eyes to read if it's in the equation form. I get the same derivative as you.

OpenStudy (anonymous):

And yet the homework site thinks my derivative is wrong for some reason. *headdesk*

OpenStudy (anonymous):

Does your homework site allow you to join all constant to that -constant term- thing, or what.

OpenStudy (anonymous):

Lemme use the equation thing to post the equation I was given for electric potential V. \[V=\frac{ \sigma }{ 2\epsilon _{0} }\sqrt{x ^{2}+R ^{2}}-x\] The problem wanted -dV/dx

OpenStudy (anonymous):

Absolutely nothing wrong w/ your answer. Hmmm....

OpenStudy (anonymous):

\[-\frac{ \sigma }{ 2\epsilon_0 }\frac{ x }{ \sqrt{x^2+R^2} }+1\] right?

OpenStudy (anonymous):

Yeah, that's right. The site calls it a partial derivative with respect to x, but dV/dx should be the same thing since the equation has no y or z in it.

OpenStudy (anonymous):

Is there a possibility for the site to wrong? I guess that's why.

OpenStudy (anonymous):

That could be. There was a question on a previous assignment that a teacher proved I did correctly but the site said my answer to was wrong.

OpenStudy (anonymous):

Ugh, finally just hit the button to see the answer they wanted. Their answer was \[\frac{ \sigma x }{ 2\epsilon _{0} }(\frac{ 1 }{ x}-\frac{ 1 }{ \sqrt{x ^{2}+R ^{2}} })\] EXACTLY my answer, just left in a weird unsimplified way. Really, homework site? DX

OpenStudy (anonymous):

You did well, that's why no substitute for the real human teacher.

OpenStudy (anonymous):

My computer programming teacher is right. Computers are dumb.

OpenStudy (anonymous):

Well, more or less.